OpenStudy (anonymous):

A college studen invests $11,000 in an account paying 8% per year compounded annually. In how many years will the amount quadruple?

OpenStudy (anonymous):

11,000(1+.08)^t = 33,000 solve for t

OpenStudy (anonymous):

I don't understand how to solve for t when it's up in a power?

OpenStudy (anonymous):

5 years

OpenStudy (anonymous):

I got 14 years...

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Well, this is what I did. 11000(1.08)^t=33000. 3=(1.08)^t log3=log(1.08) t=log(3)/log(1.08) t=14.274

OpenStudy (anonymous):

Aha I might be wrong but I got 7.53 years.

OpenStudy (anonymous):

How'd you get 7.53?!